write a rule for linear functions with graphs containing the foollowing pairs of points.
(0,3) and (6,6)

Answer by stanbon(75887)

write a rule for linear functions with graphs containing the foollowing pairs of points.
(0,3) and (6,6)
----
slope = (6-3)/(6-0) = 1/2
----
Form y = mx + b
Solve for "b":
6 = (1/2)6 + b
b = 3
----
Equation:
y = (1/2)x + 3
=================
Cheers,
Stan H.
=================